# Abelian extensions of equicharacteristic regular rings need not be Cohen-Macaulay

### v1

@@ 2. Modular invariant rings
We mention that the ring $S$ is a UFD since there is no nontrivial character $H\longrightarrow\mathbb{F}_{\!p}^{\times}$, see, for example, [Be, Corollary 3.9.3]. While $S$ is not Cohen-Macaulay, it evidently admits a small Cohen-Macaulay algebra, namely the polynomial ring $T$.
@@
Extending the example above, let $P$ be an abelian $p$-group acting on a finite rank $\mathbb{F}_{\!p}$-vector space $V$, such that the invariant ring $\mathbb{F}_{\!p}[V]^{P}$ is polynomial; this holds, for example, if the fixed subspace $V^{P}$ has codimension one, see, for example, [CW, Theorem 3.9.2]. For an integer $d\geqslant 3$, consider the $d$-fold product $G\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}P\times\dots\times P$ acting on the polynomial ring $T\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbb{F}_{\!p}[V^{\oplus d}]$. Then the invariant ring $T^{G}$ is the tensor product of $d$ copies of $\mathbb{F}_{\!p}[V]^{P}$, hence a polynomial ring. Set $H$ to be the diagonal copy of $P$ in $G$, which contains no bireflections other than the identity, since $d\geqslant 3$. One obtains a tower of rings
@@
with its usual action on the rank two vector space $\mathbb{F}_{\!p}^{2}$. Let $G$ be the product of $d$ copies of $P$ for an integer $d\geqslant 3$, and let $T$ and $H$ be as above; note that $\dim T=2d$. The fact that the invariant ring $S\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}T^{H}$ is not Cohen-Macaulay also follows from the work of Ellingsrud and Skjelbred: [ES, Corollaire 2.4] implies that $S$ does not satisfy the Serre condition $S_{3}$, while [ES, Corollaire 3.2] implies that $S$ has depth $d+2$. In particular, the Cohen-Macaulay defect of the ring $S$ is 2d(d+2)=d22d-(d+2)=d-2.

### v2

@@ 2. Modular invariant rings
We mention that the ring $S$ is a UFD since there is no nontrivial character $H\longrightarrow\mathbb{F}_{\!p}^{\times}$, see, for example, [Be, Corollary 3.9.3]. While $S$ is not Cohen-Macaulay, it evidently admits a small Cohen-Macaulay algebra, namely the polynomial ring $T$. Consequently, $T$ is not a free $S$-module.
@@
Extending the example above, let $P$ be an abelian $p$-group acting on a finite rank $\mathbb{F}_{\!p}$-vector space $V$, such that the invariant ring $\mathbb{F}_{\!p}[V]^{P}$ is polynomial; this holds, for example, if the fixed subspace $V^{P}$ has codimension one, see [CW, Theorem 3.9.2]. For an integer $d\geqslant 3$, consider the $d$-fold product $G\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}P\times\dots\times P$ acting on the polynomial ring $T\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathbb{F}_{\!p}[V^{\oplus d}]$. Then the invariant ring $T^{G}$ is the tensor product of $d$ copies of $\mathbb{F}_{\!p}[V]^{P}$, hence a polynomial ring. Set $H$ to be the diagonal copy of $P$ in $G$, which contains no bireflections other than the identity, since $d\geqslant 3$. One obtains a tower of rings
@@
with its usual action on the rank two vector space $\mathbb{F}_{\!p}^{2}$. Let $G$ be the product of $d$ copies of $P$ for an integer $d\geqslant 3$, and let $T$ and $H$ be as above; note that $\dim T=2d$. The fact that the invariant ring $S\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}T^{H}$ is not Cohen-Macaulay also follows from the work of Ellingsrud and Skjelbred: [ES, Corollaire 2.4] implies that $S$ does not satisfy the Serre condition $(S_{3})$, while [ES, Corollaire 3.2] implies that $S$ has depth $d+2$. In particular, the Cohen-Macaulay defect of the ring $S$ is 2d(d+2)=d22d-(d+2)=d-2.
While we have constructed examples with extension degree $p^{2}$, the following remains:
Question 2.4.
Let $R$ be a regular local ring of characteristic $p>0$, with fraction field $K$. Let $L$ be a Galois extension of $K$ with Galois group $\mathbb{Z}/p\mathbb{Z}$, and $S$ the integral closure of $R$ in $L$. Is $S$ a Cohen-Macaulay ring?
The analogous statement is false in mixed characteristic by [Ko, Example 2.4]. In view of Roberts’ result, as discussed in the introduction, one may ask: Let $R$ be a local UFD of characteristic $p>0$, with fraction field $K$. Let $L$ be a Galois extension of $K$ with Galois group $\mathbb{Z}/p\mathbb{Z}$, and $S$ the integral closure of $R$ in $L$. Is $S$ a free $R$-module? The answer to this is negative: the ring $T^{H}$ in Example 2.2 is a UFD, while $T$ is not a free $T^{H}$-module.
We remark that the answer to Question 2.4 is positive for $p=2$: the inclusion $R\lhook\joinrel\longrightarrow S$ splits as $R$ is regular, so $S$ is isomorphic to $R\oplus M$ for $M$ a reflexive rank one $R$-module; but then $M$ is isomorphic to $R$ since $R$ is a UFD. We can also answer Question 2.4 in the affirmative in the following special case:
Remark 2.5.
Continuing with the notation as in Question 2.4, the field extension $L/K$ is given by an Artin-Schreier polynomial $X^{p}-X-a/b\in K[X]$, with $a,b\in R$ coprime. We show that $S$ is Cohen-Macaulay if $b$ is a squarefree element of $R$. We may assume $p\geqslant 3$.
Clearing denominators, $L/K$ is also the field extension corresponding to
$F(X)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}X^{p}-b^{p-1}X-ab^{p-1}\ \in\ R[X],$
[…]
Let $\omega$ denote the image of $X$ in $A$, so $F(\omega)=0$. Let $\mathfrak{P}$ be an associated prime of $\mathfrak{C}$ in $A$. Since $A$ is $(S_{2})$ and $\mathfrak{C}$ is an $A$-dual, $\mathfrak{C}$ is $(S_{2})$, so $\mathfrak{P}$ has height one. The separability of $L/K$ implies that $F^{\prime}(\omega)=-b^{p-1}\in\mathfrak{C}$, see, for example, [SH, Theorem 12.1.1]. Thus, $\mathfrak{P}$ contains a prime factor $\pi$ of $b$ in $R$. Moreover, $\mathfrak{P}$ contains $\omega$ as $\omega^{p}=b^{p-1}(\omega+a)$. But $(\omega,\pi)A$ is a prime ideal of height one, thus it equals $\mathfrak{P}$. Hence each associated prime of $\mathfrak{C}$ is of the form $\mathfrak{P}=(\omega,\pi)A$, for some prime factor $\pi$ of $b$ in $R$. We next compute $\mathfrak{C}_{\mathfrak{P}}$.
As $b$ is squarefree, we may rewrite $F=X^{p}-b^{p-1}(X+a)$ as $F=X^{p}-\pi^{p-1}\lambda$ for some unit $\lambda\in R[X]_{(X,\pi)}$. The element $\pi/\omega$ is integral over $A_{\mathfrak{P}}$ since $(\pi/\omega)^{p-1}=\lambda^{-1}\omega$, so
$A_{\mathfrak{P}}[\pi/\omega]\ \subseteq\ S_{\mathfrak{P}}.$
We claim that equality holds above. Each maximal ideal of $A_{\mathfrak{P}}[\pi/\omega]$ contains $\omega$, $\pi$, and also $\pi/\omega$, using once again that $(\pi/\omega)^{p-1}=\lambda^{-1}\omega$. But $(\pi,\omega,\pi/\omega)A_{\mathfrak{P}}[\pi/\omega]$ is a maximal ideal, so $A_{\mathfrak{P}}[\pi/\omega]$ is local, moreover with maximal ideal generated by $\pi/\omega$. It follows that $A_{\mathfrak{P}}[\pi/\omega]$ is a regular local ring, hence normal, proving the claim. Note that $S_{\mathfrak{P}}$ is a discrete valuation ring with uniformizer $\pi/\omega$, while the maximal ideal $(\omega,\pi)A_{\mathfrak{P}}$ of the subring $A_{\mathfrak{P}}$ is generated by
$(\pi/\omega)^{p-1}=\lambda^{-1}\omega\quad\text{ and }\quad(\pi/\omega)^{p}=\lambda^{-1}\pi.$
Upon completion, the inclusion $A_{\mathfrak{P}}\subset S_{\mathfrak{P}}$ takes the form
$\mathbb{F}\llbracket(\pi/\omega)^{p-1},\,(\pi/\omega)^{p}\rrbracket\ \subset\ \mathbb{F}\llbracket\pi/\omega\rrbracket,$
for $\mathbb{F}$ a coefficient field, so the calculation of the conductor is essentially the calculation of the Frobenius number of the semigroup generated by p1p-1 and $p$, and this is $p^{2}-3p+1$. It follows that $\mathfrak{C}_{\mathfrak{P}}=\mathfrak{P}^{p-2}A_{\mathfrak{P}}$, and, in turn, that $\mathfrak{C}A_{\mathfrak{P}}\cap A$ is the symbolic power $\mathfrak{P}^{(p-2)}$.
As $\mathfrak{C}$ is unmixed, we obtain the primary decomposition $\mathfrak{C}=\bigcap_{i}(\omega,\pi_{i})^{(p-2)}$, where $\pi_{i}$ ranges over the prime factors of $b$ in $R$. Note that $F\in(X,\pi_{i})^{p-2}$ for each $i$, and thus,
$A/\mathfrak{C}\ \cong\ \frac{R[X]}{\bigcap_{i}(X,\pi_{i})^{(p-2)}}\ =\ \frac{R[X]}{\bigcap_{i}(X,\pi_{i})^{p-2}},$
where the equality holds as ordinary and symbolic powers coincide for ideals generated by a regular sequence. Moreover, in the regular ring $R[X]$, it is readily seen that
$\bigcap\nolimits_{i}(X,\pi_{i})^{p-2}\ =\ (X,b)^{p-2}$
since $b$ is squarefree, with the elements $\pi_{i}$ as its prime factors. Since $(X,b)^{p-2}$ is a perfect ideal of $R[X]$, we conclude that the ring $A/\mathfrak{C}\cong R[X]/(X,b)^{p-2}$ is Cohen-Macaulay.